### v2

@@ Abstract
Vision-language models (VLMs) integrate visual and textual information, enabling a wide range of applications such as image captioning and visual question answering, making them crucial for modern AI systems. However, their high computational demands pose challenges for real-time applications. This has led to a growing focus on developing efficient vision-language models. In this survey, we review key techniques for optimizing VLMs on edge and resource-constrained devices. We also explore compact VLM architectures, frameworks and provide detailed insights into the performance-memory trade-offs of efficient VLMs. Furthermore, we establish a GitHub repository at MPSC-GitHub to compile all surveyed papers, which we will actively update. Our objective is to foster deeper research in this area.
@@ 1 Introduction
The rest of the survey is organized as follows. Section 2 covers pre-deployment techniques for efficient VLMs, while Sections 3 and 4 detail fine-tuning and runtime optimization methods. Section 5 explores distributed VLMs. Section 6 lists state-of-the-art, efficient VLMs and also details various VLM frameworks and libraries. Section 7 provides insights into the accuracy vs. efficiency trade-offs for VLMs. Section 8 outlines application avenues, Section 9 discusses challenges and future directions, and Section 10 concludes the paper.
We also create a GitHub repository to compile the papers featured in this survey: MPSC-GitHub. It will be actively maintained with updates on emerging research.
@@ 2 Pre-deployment Techniques
Pre-deployment techniques refer to optimization methods applied to models before deployment to improve their efficiency while maintaining accuracy. In our survey, we focus on techniques such as quantization, low-rank approximation, pruning, and knowledge distillation. These methods are particularly suited for edge devices as they reduce computational and memory requirements. The most commonly used benchmarked datasets for evaluating these methods include ScienceQA [120], VQA-v2 [54], MMMU [205], SEED [91], and OCRBench [112]. The primary evaluation metrics are accuracy, top-1 accuracy, and the harmonic mean, as employed in Q-VLM [174], MBQ [95], MoPE-CLIP [104], and PromptKD [100]. Inference latency and memory consumption are commonly used to assess the efficiency of these approaches. Training is typically conducted on high-performance GPUs such as the RTX 4090, V100, and A100.
@@ 2.1 Quantization
$q(v)=\text{round}(\eta\cdot v).$ (4)
@@
$q(v)=\text{round}(\eta\cdot v+\gamma).$ (5)
@@
$\eta=\frac{\theta_{\text{max}}-\theta_{\text{min}}}{2^{n-1}},\quad\gamma=\text{round}(-\eta\cdot\theta_{\text{min}}).$ (6)
@@ 2.1.1 Post-Training Quantization (PTQ)
PTQ updates the weights and activations after training using a small dataset, employing either symmetric or asymmetric methods. It is simpler for weights due to direct access to weight tensors but more challenging for activations as they require measurement with input data. This requires innovative techniques to balance efficiency and accuracy, especially when handling the complexities of activation quantization. [174] propose Q-VLM, a PTQ framework for large vision-language models (LVLMs) that focuses on minimizing cross-layer dependency errors using activation entropy as a proxy. This method partitions the model into blocks for efficient quantization and employs visual encoder optimization to reduce search space while maintaining accuracy. It achieves significant memory compression and faster inference speeds without performance degradation. [95] introduces Modality-Balanced Quantization (MBQ) for LVLMs that accounts for sensitivity differences between vision and language tokens. By balancing reconstruction loss during calibration, MBQ improves accuracy and supports weight-only and weight-activation coquantization, achieving up to 1.4× decoding speedup with a custom W3 kernel. NoisyQuant [114] reshapes activation distributions using uniform noisy bias to reduce quantization errors in vision transformers and can be adapted to VLMs for efficient activation quantization. To mitigate the performance drop problem in PTQ, P4Q [163] introduces learnable prompts and a lightweight low-bit adapter to realign image and text feature distributions. This approach enhances recognition performance and achieves results comparable to full-precision models while maintaining computational efficiency. Some PTQ methods are specifically designed for vision and language models independently; however, they can also be applied to vision and language components within VLMs. Therefore, it is essential to consider these methods when discussing PTQ strategies for VLMs. [116] propose a PTQ method tailored for vision transformers, introducing similarity- and ranking-aware objectives to preserve attention behavior without fine-tuning. They also use mixed-precision quantization based on nuclear norms, achieving near-full precision accuracy (81. 29% top-1 on ImageNet with DeiT-B) while significantly reducing model size and latency. [107] present FQ-ViT as a post-training method that enables fully quantized vision transformers by addressing LayerNorm and Softmax issues using Power-of-Two Factor (PTF) and Log-Int-Softmax (LIS). This approach achieves 84.89 % top-1 accuracy with ViT-L on ImageNet. [215] introduce SageAttention, a plug-and-play PTQ method that quantizes the attention module in Transformers. It achieves INT8 quantization of attention while preserving accuracy through innovations like smoothing the K matrix and using FP16 accumulators, resulting in up to 2.1× faster inference compared to FlashAttention2 [25] across various models. Building on this, SageAttention2 [211] further accelerates attention using INT4 per-thread quantization alongside precision-enhancing techniques such as Q-smoothing and two-level accumulation. It achieves up to 3× faster inference than FlashAttention2 while maintaining end-to-end accuracy across language, image, and video generation models. [204] propose twin uniform quantization for vision transformers to handle the special distributions of activations and a Hessian-guided metric to improve scaling factor selection. This can enhance the efficiency of transformer-based vision encoders in VLMs while maintaining near-lossless accuracy. Similarly, [121] develop PTQ4SAM, a quantization framework for the Segment Anything Model (SAM), which incorporates a Bimodal Integration strategy to address challenging activation distributions and Adaptive Granularity Quantization to improve softmax quantization. PTQ techniques such as LRQuant and BiLLM offer promising solutions for compressing language components in VLMs. LRQuant [217] enhances learnability and generalization through dynamic smoothing and test-time adaptation, while BiLLM [65] improves and extends ultra-low-bit quantization to minimize memory and computation costs. SmoothQuant [183] enables training-free quantization for LLMs by smoothing activation outliers and balancing quantization difficulty via per-channel scaling, achieving W8A8 quantization with up to 1.56× speedup and minimal accuracy loss across models up to 530B on a single node.
@@ 2.1.2 Quantization-Aware Training (QAT)
QAT occurs during training itself, allowing the model to fine-tune with quantized weights and achieve better performance than PTQ. Fake quantization is applied in the forward pass, making the model adapt to quantization effects while retaining the full precision version. Very few studies have directly investigated this approach in the context of VLMs. [184] introduce QSLAW, a quantization-aware scale learning approach for VLMs, optimizing multimodal large language models (MLLMs) by learning group-wise scale factors to mitigate quantization errors. Additionally, it uses a modality-aware warmup approach to enhance VL instruction tuning efficiency while  preserving linguistic knowledge. Most studies have focused on either the vision or language component, but surveying these methods is valuable as they can be applied separately. [66] identify quantization variation as a key challenge in transformer QAT and propose a variation-aware scheme combining module-dependent scaling, oscillation-aware regularization, and knowledge distillation to achieve state-of-the-art low-bit performance with improved training efficiency. Fixed-point ViT [101] introduces an enhanced QAT strategy by combining mean-square error-based initialization with learned step size quantization, achieving nearly-lossless 8-bit quantization. Q-ViT [102] proposes a fully differentiable QAT framework for Vision Transformers, optimizing both quantization scales and bit-widths dynamically. It employs head-wise bit-width allocation and a switchable scale mechanism to enhance quantization robustness, achieving 3-bit precision with minimal accuracy loss. GPUSQ-ViT [201] employs sparse distillation-aware QAT to optimize Vision Transformers for GPU efficiency. By integrating feature-based knowledge distillation with quantization scaling, it minimizes accuracy degradation while achieving substantial reductions in model size. PackQViT [32] introduces activation-aware sub-8-bit QAT with log2 and outlier-aware quantization, achieving real-time ViT inference and superior accuracy on mobile devices. LLM-QAT [115] leverages self-generated data from the pre-trained model for knowledge distillation, enabling QAT without access to the original dataset. It quantizes weights, activations, and the KV cache and significantly improves performance over PTQ methods. [20] employs QAT (EfficientQAT) to optimize LLMs by introducing block-wise training of all parameters (Block-AP) and end-to-end training of quantization parameters (E2E-QP). This approach reduces memory overhead and training time while maintaining high accuracy in low-bit quantization scenarios. QA-LoRA [191] applies group-wise quantization alongside low-rank adaptation. This balances the flexibility between quantization and adaptation, enabling efficient training and deployment without the need for post-training quantization. LR-QAT [12] is a lightweight, memory-efficient QAT method for LLMs that combines low-rank adaptation with fixed-point downcasting and checkpointing. It achieves full-model QAT performance with minimal overhead and also supports various quantization settings.
@@ 2.2 Low-rank Approximation
Low-rank approximation has been widely used for language models and has recently been applied to multimodal models as well. [158] introduce LoRA-Sparse, a low-rank approximation method for sparse attention in LLMs and multimodal models like LLaVA. LoRA-Sparse employs order-mimic training to approximate full attention effectively, demonstrating benefits across NLP and multimodal benchmarks. SeTAR [98] proposes a training-free selective low-rank approximation method for Out of Distribution (OOD) detection in CLIP-based models, enhancing efficiency and accuracy without requiring fine-tuning. SeTAR+FT [98] extends this by integrating fine-tuning for further performance gains, achieving state-of-the-art results on ImageNet1K and Pascal-VOC benchmarks. PELA [59] incorporates an intermediate pre-training stage with low-rank approximation to compress pre-trained models. It enables efficient fine-tuning by distilling features and regularizing weight perturbations to improve scalability for downstream tasks.
@@ 2.3 Pruning
Parameter pruning involves removing redundant or less significant model weights. A simple yet effective pruning method is to remove model weights that fall below a predetermined threshold. This thresholding approach is often referred to as magnitude pruning and serves as the basis for many iterative pruning strategies. However, more advanced techniques exist, such as gradient-based pruning, which evaluates weight importance based on sensitivity to loss function, and Hessian-based pruning that uses second-order information to identify less impactful weights. Parameter pruning reduces the memory footprint and computational load of large networks, making them more efficient for resource-constrained devices. By eliminating redundant weights, pruning enables faster inference, lower power consumption, and efficient deployment on edge devices. Pruning is categorized into Structured and Unstructured Pruning. A schematic representation for the same is shown in figure 5.
@@ 2.4 Knowledge Distillation (KD)
KD is a model compression technique in which a teacher network transfers knowledge to a smaller student network. The teacher network is typically a large neural network with many parameters. Knowledge can reside in internal layers, activations, or even soft labels 6. KD enables deployment of a smaller high-performance model on edge devices. Several studies have used KD in VLMs to enhance efficiency. [100] introduce PromptKD, a novel approach that distills knowledge from a pretrained CLIP teacher model to a lightweight student using prompt-based imitation, enabling effective adaptation to domain-specific tasks without requiring labeled data. PromptKD additionally allows fast inference by leveraging pre-stored text features. VLMs still struggle with understanding spatial and contextual relationships. To address this SF-CLIP [149] employ masked knowledge distillation from pre-trained vision and language models, ensuring better spatial feature learning and multilingual retrieval while preserving training efficiency. [39] propose a distillation approach for compressing large visual-linguistic models by aligning student and teacher model representations. They ensure consistency in visual tokens using the same region proposals and introduce loss functions to match attention distributions and hidden states, improving efficiency without sacrificing performance. [113] develop KD-VLP, an object-aware framework that embeds object-level knowledge into multimodal learning. By leveraging knowledge distillation from object features and semantic labels, their method improves cross-modal alignment and enhances downstream task performance without requiring external object detectors during inference. VLM-based detectors struggle with domain adaptation due to reliance on biased object proposals from the target domain, making direct knowledge distillation ineffective. To solve this issue, VLDadaptor [79] leverage domain-mixed contrastive knowledge distillation and domain-mixed consistency distillation to align category-level features in the visual-language embedding space. Even the recent breakthrough, DeepSeek-R1 [26], demonstrates that distilling its outputs to a student model surpasses previous non-reasoning models.
@@ 2.5 Other Methods
Recent pre-deployment work has focused on developing mixture of experts (MoE) and adaptive attention mechanisms for resource-constrained devices. Mixture of Experts enhances model efficiency by dynamically selecting specialized subnetworks or "experts," for different inputs. MoE facilitates accelerated inference relative to a model with the same number of parameters. Figure 7 shows the functioning of MoE with a gating mechanism to pass information to multiple experts. Load balancing is essential during training to avoid bias toward specific experts [178]. Recent research leverages MoE to enhance efficiency in VLMs. [202] introduce a parameter-efficient continuous learning framework for VLMs by integrating MoE adapters. MoE facilitates zero-shot capabilities through the dynamic selection of task-specific adapters. [153] illustrate the importance of MoE in scaling VLMs while balancing the trade-off between computational resources and performance. Med-MoE [70] leverages domain-specific experts and a meta-expert to efficiently handle generative and discriminative medical tasks while reducing the number of parameters during inference. Another research focus is optimizing attention mechanisms to maintain performance while reducing FLOPs. Adaptive attention [118] has emerged as a promising approach that  selectively attends to tokens based on context, as compared to previous mechanisms that attended to all tokens [173]. The above described evaluation metrics and datasets are used for this as well.
@@ 3 Efficient Fine-tuning
Efficient fine-tuning refers to optimizing pre-trained models by updating only a small subset of parameters. Fine-tuning eliminates the need for full retraining, thereby reducing memory consumption. Fine-tuning techniques include both parameter-efficient and memory-efficient methods. The most commonly used datasets remain largely consistent with previous works, with notable additions such as MSCOCO [105], MSR-VTT [189], and ImageNet [27]. Evaluation metrics include recall@k [68], harmonic mean (HM), few-shot accuracy [197], and domain generalization [117]. Memory consumption is frequently reported to assess efficiency. Experiments are conducted on V100 and RTX 3090 GPUs.
@@ 3.1.1 Low-Rank Adaptation (LoRA)
Standard fine-tuning requires updating all model parameters, which is not only time-intensive but also requires substantial storage. In contrast, LoRA uses low-rank decomposition to approximate the weight update matrix (refer figure 8). Instead of altering the entire weight matrix $W\in\mathbb{R}^{m\times n}$, LoRA parameterizes the update matrix $\Delta{W}$ as the product of two lower-dimensional matrices, capturing essential task-specific adaptations. LoRA decomposes the update matrix $\Delta W$ as:
$\Delta W=PQ^{T},\quad\text{where }P\in\mathbb{R}^{m\times r},\quad Q\in\mathbb{R}^{n\times r}.$ (9)
The updated weight matrix is given by:
$W^{*}=W+\Delta W=W+PQ^{T}.$ (10)
With VLMs gaining significant attention, studies have explored applying LoRA to enhance their efficiency. [9] propose an adapter ensemble strategy to efficiently fine-tune large-scale VLMs while integrating LoRA to mitigate the additional parameter burden introduced by the ensemble process. By including LoRA, they aim to reduce memory and compute overhead while maintaining competitive performance, particularly in vision-language retrieval tasks. CLIP-LoRA [207] reduces training overhead while maintaining strong performance across 11 datasets, demonstrating that LoRA can outperform prompt tuning and adapter-based approaches in few-shot scenarios without requiring extensive hyperparameter tuning. AdvLoRA [68] builds on LoRA by introducing a parameter-efficient adversarial adaptation method for VLMs, incorporating novel reparameterization, parameter clustering, and adaptive updates to enhance both robustness and efficiency under adversarial conditions. Comp-LoRA extends LoRA [181] for few-shot fine-tuning of VLMs by optimizing low-rank updates in the complementary subspace, effectively mitigating catastrophic forgetting and preserving the model’s original vision-language alignment capabilities. Block-LoRA [221] introduces a block matrix-based approach that partitions LoRA’s low-rank matrices into sub-matrices with shared components. This leads to a reduction in the number of training parameters and simplifies computations for CLIP-based few-shot learning while retaining competitive performance. CoDyRA [117] proposes a dynamic rank-selective LoRA for continual learning in VLMs, adaptively adjusting the rank of low-rank updates based on task importance. It preserves pre-trained knowledge while enabling efficient adaptation without additional inference overhead. LoRA-TTT [87] applies low-rank adaptation to the image encoder of VLMs during test time, enabling efficient domain adaptation without tuning text prompts, thereby improving generalization with minimal memory and runtime overhead. MTLoRA [3] enables efficient multi-task adaptation by using task-agnostic and task-specific low-rank modules, achieving higher accuracy with 3.6× fewer trainable parameters. Serial LoRA [219] uses a shared low-rank matrix across attention heads in vision transformers, reducing parameters to 1/4 of standard LoRA while maintaining comparable performance.

### v3

@@ Abstract
Vision-language models (VLMs) integrate visual and textual information, enabling a wide range of applications such as image captioning and visual question answering, making them crucial for modern AI systems. However, their high computational demands pose challenges for real-time applications. This has led to a growing focus on developing efficient vision-language models. In this survey, we review key techniques for optimizing VLMs on edge and resource-constrained devices. We also explore compact VLM architectures, frameworks and provide detailed insights into the performance-memory trade-offs of efficient VLMs. Furthermore, we establish a GitHub repository at https://github.com/MPSC-UMBC/Efficient-Vision-Language-Models-A-Survey to compile all surveyed papers, which we will actively update. Our objective is to foster deeper research in this area.
@@ 1 Introduction
The rest of the survey is organized as follows. Section 2 covers pre-deployment techniques for efficient VLMs, while Sections 3 and 4 detail fine-tuning and runtime optimization methods. Section 5 explores distributed VLMs. Section 6 lists state-of-the-art, efficient VLMs and also details various VLM frameworks and libraries. Section 7 provides insights into the accuracy vs. efficiency trade-offs for VLMs. Section 8 outlines application avenues, Section 9 discusses challenges and future directions, and Section LABEL:conclusionfuturedirections concludes the paper.
We also create a GitHub repository to compile the papers featured in this survey: https://github.com/MPSC-UMBC/Efficient-Vision-Language-Models-A-Survey. It will be actively maintained with updates on emerging research.
@@ 2 Pre-deployment Techniques
Pre-deployment techniques refer to optimization methods applied to models before deployment to improve their efficiency while maintaining accuracy. In our survey, we focus on techniques such as quantization, low-rank approximation, pruning, and knowledge distillation. These methods are particularly suited for edge devices as they reduce computational and memory requirements.
@@ 2.1 Quantization
$q(v)=\text{round}(\eta\cdot v)$ (4)
@@
$q(v)=\text{round}(\eta\cdot v+\gamma)$ (5)
@@
$\eta=\frac{\theta_{\text{max}}-\theta_{\text{min}}}{2^{n-1}},\quad\gamma=\text{round}(-\eta\cdot\theta_{\text{min}})$ (6)
@@ 2.1.1 Post-Training Quantization (PTQ)
PTQ updates the weights and activations after training using a small dataset, employing either symmetric or asymmetric methods. It is simpler for weights due to direct access to weight tensors but more challenging for activations as they require measurement with input data. This requires innovative techniques to balance efficiency and accuracy, especially when handling the complexities of activation quantization. [143] propose Q-VLM, a PTQ framework for large vision-language models (LVLMs) that focuses on minimizing cross-layer dependency errors using activation entropy as a proxy. This method partitions the model into blocks for efficient quantization and employs visual encoder optimization to reduce search space while maintaining accuracy. It achieves significant memory compression and faster inference speeds without performance degradation. [78] introduce Modality-Balanced Quantization (MBQ) for LVLMs that accounts for sensitivity differences between vision and language tokens. By balancing reconstruction loss during calibration, MBQ improves accuracy and supports weight-only and weight-activation coquantization, achieving up to 1.4× decoding speedup with a custom W3 kernel. NoisyQuant [92] reshapes activation distributions using uniform noisy bias to reduce quantization errors in vision transformers and can be adapted to VLMs for efficient activation quantization. To mitigate the performance drop problem in PTQ, P4Q [133] introduces learnable prompts and a lightweight low-bit adapter to realign image and text feature distributions. This approach enhances recognition performance and achieves results comparable to full-precision models while maintaining computational efficiency. Some PTQ methods are specifically designed for vision and language models independently; however, they can also be applied to vision and language components within VLMs. Therefore, it is essential to consider these methods when discussing PTQ strategies for VLMs. [170] propose twin uniform quantization for vision transformers to handle the special distributions of activations and a Hessian-guided metric to improve scaling factor selection. This can enhance the efficiency of transformer-based vision encoders in VLMs while maintaining near-lossless accuracy. Similarly, [95] develop PTQ4SAM, a quantization framework for the Segment Anything Model (SAM), which incorporates a Bimodal Integration strategy to address challenging activation distributions and Adaptive Granularity Quantization to improve softmax quantization. PTQ techniques such as LRQuant and BiLLM offer promising solutions for compressing language components in VLMs. LRQuant [179] enhances learnability and generalization through dynamic smoothing and test-time adaptation, while BiLLM [55] improves and extends ultra-low-bit quantization to minimize memory and computation costs.
@@ 2.1.2 Quantization-Aware Training (QAT)
QAT occurs during training itself, allowing the model to fine-tune with quantized weights and achieve better performance than PTQ. Fake quantization is applied in the forward pass, making the model adapt to quantization effects while retaining the full precision version. Very few studies have directly investigated this approach in the context of VLMs. [151] introduce QSLAW, a quantization-aware scale learning approach for VLMs, optimizing multimodal large language models (MLLMs) by learning group-wise scale factors to mitigate quantization errors. Additionally, it uses a modality-aware warmup approach to enhance VL instruction tuning efficiency while  preserving linguistic knowledge. Most studies have focused on either the vision or language component, but surveying these methods is valuable as they can be integrated. Q-ViT [84] proposes a fully differentiable QAT framework for Vision Transformers, optimizing both quantization scales and bit-widths dynamically. It employs head-wise bit-width allocation and a switchable scale mechanism to enhance quantization robustness, achieving 3-bit precision with minimal accuracy loss. GPUSQ-ViT [166] employs sparse distillation-aware QAT to optimize Vision Transformers for GPU efficiency. By integrating feature-based knowledge distillation with quantization scaling, it minimizes accuracy degradation while achieving substantial reductions in model size. LLM-QAT [93] leverages self-generated data from the pre-trained model for knowledge distillation, enabling QAT without access to the original dataset. It quantizes weights, activations, and the KV cache and significantly improves performance over PTQ methods. [15] employs QAT (EfficientQAT) to optimize LLMs by introducing block-wise training of all parameters (Block-AP) and end-to-end training of quantization parameters (E2E-QP). This approach reduces memory overhead and training time while maintaining high accuracy in low-bit quantization scenarios.
@@ 2.2 Low-rank Approximation
Low-rank approximation has been widely used for language models and has recently been applied to multimodal models as well. [128] introduce LoRA-Sparse, a low-rank approximation method for sparse attention in LLMs and multimodal models like LLaVA. LoRA-Sparse employs order-mimic training to approximate full attention effectively, demonstrating benefits across NLP and multimodal benchmarks. SeTAR [81] proposes a training-free selective low-rank approximation method for OOD detection in CLIP-based models, enhancing efficiency and accuracy without requiring fine-tuning. SeTAR+FT [81] extends this by integrating fine-tuning for further performance gains, achieving state-of-the-art results on ImageNet1K and Pascal-VOC benchmarks. PELA [49] incorporates an intermediate pre-training stage with low-rank approximation to compress pre-trained models. It enables efficient fine-tuning by distilling features and regularizing weight perturbations to improve scalability for downstream tasks.
@@ 2.3 Pruning
Parameter pruning involves removing redundant or less significant model weights. A simple yet effective pruning method is to remove model weights that fall below a predetermined threshold. However, more advanced techniques exist, such as gradient-based pruning, which evaluates weight importance based on sensitivity to loss function, and Hessian-based pruning that uses second-order information to identify less impactful weights. Parameter pruning reduces the memory footprint and computational load of large networks, making them more efficient for resource-constrained devices. By eliminating redundant weights, pruning enables faster inference, lower power consumption, and efficient deployment on edge devices. Pruning is categorized into Structured and Unstructured Pruning. A schematic representation for the same is shown in figure 5.
@@ 2.4 Knowledge Distillation (KD)
KD is a model compression technique in which a teacher network transfers knowledge to a smaller student network. The teacher network is typically a large neural network with many parameters. Knowledge can reside in internal layers, activations, or even soft labels 6. KD enables deployment of a smaller high-performance model on edge devices. Several studies have used KD in VLMs to enhance efficiency. [83] introduce PromptKD, a novel approach that distills knowledge from a pretrained CLIP teacher model to a lightweight student using prompt-based imitation, enabling effective adaptation to domain-specific tasks without requiring labeled data. PromptKD additionally allows fast inference by leveraging pre-stored text features. VLMs still struggle with understanding spatial and contextual relationships. To address this SF-CLIP [120] employ masked knowledge distillation from pre-trained vision and language models, ensuring better spatial feature learning and multilingual retrieval while preserving training efficiency. [31] propose a distillation approach for compressing large visual-linguistic models by aligning student and teacher model representations. They ensure consistency in visual tokens using the same region proposals and introduce loss functions to match attention distributions and hidden states, improving efficiency without sacrificing performance. [91] develop KD-VLP, an object-aware framework that embeds object-level knowledge into multimodal learning. By leveraging knowledge distillation from object features and semantic labels, their method improves cross-modal alignment and enhances downstream task performance without requiring external object detectors during inference. VLM-based detectors struggle with domain adaptation due to reliance on biased object proposals from the target domain, making direct knowledge distillation ineffective. To solve this issue, VLDadaptor [67] leverage domain-mixed contrastive knowledge distillation and domain-mixed consistency distillation to align category-level features in the visual-language embedding space. Even the recent breakthrough, DeepSeek-R1 [20], demonstrates that distilling its outputs to a student model (DeepSeek-R1-7B) surpasses previous non-reasoning models.
@@ 2.5 Other Methods
Recent pre-deployment work has focused on developing mixture of experts (MoE) and adaptive attention mechanisms for resource-constrained devices. Mixture of Experts enhances model efficiency by dynamically selecting specialized subnetworks or "experts," for different inputs. MoE facilitates accelerated inference relative to a model with the same number of parameters. Figure 7 shows the functioning of MoE with a gating mechanism to pass information to multiple experts. Load balancing is essential during training to avoid bias toward specific experts [147]. Recent research leverages MoE to enhance efficiency in VLMs. [167] introduce a parameter-efficient continuous learning framework for VLMs by integrating MoE adapters. MoE facilitates zero-shot capabilities through the dynamic selection of task-specific adapters. [124] illustrate the importance of MoE in scaling VLMs while balancing the trade-off between computational resources and performance. Med-MoE [57] leverages domain-specific experts and a meta-expert to efficiently handle generative and discriminative medical tasks while reducing the number of parameters during inference. Another research focus is optimizing attention mechanisms to maintain performance while reducing FLOPs. Adaptive attention [94] has emerged as a promising approach that  selectively attends to tokens based on context, as compared to previous mechanisms that attended to all tokens [142].
@@ 3 Efficient Fine-tuning
Efficient fine-tuning refers to optimizing pre-trained models by updating only a small subset of parameters. Fine-tuning eliminates the need for full retraining, thereby reducing memory consumption. Fine-tuning techniques include both parameter-efficient and memory-efficient methods.
@@ 3.1.1 Low-Rank Adaptation (LoRA)
Standard fine-tuning requires updating all model parameters, which is not only time-intensive but also requires substantial storage. In contrast, LoRA uses low-rank decomposition to approximate the weight update matrix (refer figure 8). Instead of altering the entire weight matrix $W\in\mathbb{R}^{m\times n}$, LoRA parameterizes the update matrix $\Delta{W}$ as the product of two lower-dimensional matrices, capturing essential task-specific adaptations while minimizing storage and computational costs. LoRA decomposes the update matrix $\Delta W$ as:
$\Delta W=PQ^{T},\quad\text{where }P\in\mathbb{R}^{m\times r},\quad Q\in\mathbb{R}^{n\times r}$ (9)
The updated weight matrix is given by:
$W^{*}=W+\Delta W=W+PQ^{T}$ (10)
@@
With VLMs gaining significant attention, some studies have explored applying LoRA to enhance their efficiency. [7] propose an adapter ensemble strategy to efficiently fine-tune large-scale VLMs while integrating LoRA to mitigate the additional parameter burden introduced by the ensemble process. By including LoRA, they aim to reduce memory and compute overhead while maintaining competitive performance, particularly in vision-language retrieval tasks. CLIP-LoRA [172] reduces training overhead while maintaining strong performance across 11 datasets, demonstrating that LoRA can outperform prompt tuning and adapter-based approaches in few-shot scenarios without requiring extensive hyperparameter tuning. [163, 3, 186] have focused on vision models, while [58, 169] have explored language models; however, limited research has been conducted on multimodal settings like VLMs.